Suppose a firm has a net income of $600, cash dividends of $200, total assets of $4,500 and liabilities of $2,000. What is the internal growth rate?

4.4 percent

8.0 percent

8.9 percent

13.3 percent

16.0 percent
Answer #1

ROA=Net income/Total assets

=(600/4500)=0.133(Approx)

Dividend payout ratio=Dividend/Net income

=(200/600)=0.333

Retention ratio=1-Dividend payout ratio

=(1-0.333)=0.667

Hence internal growth rate=(ROA*Retention rate)

=(0.133*0.667)

=8.9%(Approx).